Tran Company, which uses the high-low method to analyze cost behavior, has determined that machine hours best predict the comDrop down boxes options and data table provided

Will give rating upon completion~
Answer #1

Solution 1:

Change in cost / change in volume = variable cost

= (4414 -3711) / (1440 -1060) = variable cost

= 703 / 380 = 1.85

Variable utility cost per machine hour = $1.85

Solution 2:

Total operating cost - total variable cost = Fixd cost

$4414 - (1.85*1440) = fixed cost

Fixed cost = $1750

Solution 3:

Total cost when 1250 machine hours used = $1750 + $1.85*1250 = $4063